# TinySQL: A Progressive Text-to-SQL Dataset for Mechanistic Interpretability Research

### v1

@@ 5 Discussion
5.1 Differences in MI techniques
Our research demonstrated that different mechanistic interpretability methods can yield varying results. We observed these differences not only between different approaches like SAEs and EAP, but also within individual methods. For instance, our activation patching experiments on a specific subset CS1 prompts (Section 4.2) produced different results, even when the prompts had largely similar structures. Furthermore, researchers must navigate many nuanced decisions that could change results, such as choosing between zero or mean ablation (Miller et al., 2024), or whether to patch a corrupted output into a clean run or vice versa (Heimersheim and Nanda, 2024).
This variability in results represents a fundamental challenge in mechanistic interpretability research. Hence, we believe that TinySQL’s controlled progression of tasks and systematic structure makes it an ideal platform for investigating these methodological differences and developing more reliable interpretability techniques.
5.2 Improving Synthetic Datasets through MI Analysis
Our initial dataset used newlines to make the SQL query easier to read, but the trained models ended up relying on these newlines to process the field names, leading to poor generalization on differently formatted queries. By removing these formatting artifacts, we stopped the model from exploiting unintended signals and created a cleaner dataset that better tested SQL understanding.
Furthermore, we found that models trained on CS1 to CS3 were predicting the field name based solely on the user query, and not the table schema. This enabled an iterative refinement process where we could identify problematic patterns in model behavior, trace them back to specific dataset characteristics, and make targeted improvements to the synthetic data. This approach helped ensure our datasets actually test the capabilities we intend to measure, rather than allowing models to succeed through undesired shortcuts. Hence, we believe that MI could help improve synthetic datasets on programmatic tasks in this manner.
@@ 6 Conclusion
We present TinySQL, a dataset with increasingly complex SQL command subsets, as well as trained models of differing sizes. Our analysis of multiple interpretability methods (EAP, SAE, and activation patching) reveals both consistent patterns and limitations in interpreting models, highlighting areas where current interpretability techniques need improvement while providing a testbed for improving the robustness of interpretability techniques.
@@
Time constraints meant we did not include planned more-complex SQL subsets including SQL joins, where clauses and multiple tables. Our findings may not generalize to all SQL dialects or database systems. Also, our mechanistic understanding of the model algorithm is incomplete, and so also is our ability to compare algorithms between models.
While we were to leverage our MI results to improve our training data, we were not able to use our MI results to improve model performance post-training. Given time constraints, we did not run all our analyses on all models, e.g. while we trained SAEs on all of BM1 through BM3, we only studied these on BM1.
@@ C.1 Dataset Statistics for CSx_Syn
C.2 Dataset Generation Methodology: CS1_Nat
@@ Appendix D Trained Models
Appendix E Activation Patching Results
Across prompts across all classes, we find that the activation patching results differ quite significantly from each other, even for prompts in the same class. We provide a sample of 4 examples in Figure 9 in App.E, and we note that this variation persists throughout all examples.
@@ Appendix E Activation Patching Results
Appendix F Circuit Identification
F.1 Edge Attribution Patching (EAP) Results
@@
F.1.1 Features used per dataset
@@ F.1.2 EAP results
F.2 Ablation Study Results
@@
F.3 Multi-Layer Perceptron (MLP) Ablations
@@
F.4 Full-Layer Ablations
@@ Appendix G SAE analysis
We show now the SAE features most involved in filling out the table name tokens, in particular for BM1_CS1_Syn, BM1_CS2_Syn and BM1_CS3_Syn. See Figures 10, 11 and 12 for the top activating SAE features by average magnitude.
@@
We can map these to the actual attention outputs, see for instance Figures 13(a), 13(b) and 13(c). We average over these to determine the influential attention heads.
@@
Figure 13: Attention outputs by head for selected SAE features across models and layers.
@@
Figure 17: Ablation results showing accuracy trends as more attention heads remain unablated. (Top-left) BM1_CS1_Syn with CS1 data, (Top-right) BM1_CS3_Syn with CS1 data, (Bottom) BM1_CS3_Syn with CS2 data.

### v2

@@ 7 Discussion
7.1 Ablation Validates Robustness–Granularity Tradeoff
Our ablation results reveal that zero-ablation recovers performance significantly better than random circuits, making it more reliable for identifying robust, task-relevant components. In contrast, mean-ablation yields smaller circuits but recovers performance similar to random baselines, suggesting weaker causal validity. Thus, zero-ablation is better for robustness, while mean-ablation may be useful for identifying compact, interpretable circuits when some loss in reliability is acceptable. Notably, across prompts within all classes, activation patching results vary significantly, even among prompts in the same class, highlighting the inherent variability in the circuits identified (see Appendix F for examples).
7.2 Distributed Computation in SQL Generation
LogitLens analysis shows that SQL generation unfolds compositionally: early layers capture intent, middle layers resolve schema references, and later layers assemble the final query. However, EAP often reveals fragmented mechanisms scattered across layers, especially in larger models, highlighting the difficulty of isolating clean circuits and the limits of current interpretability tools at scale. This fragmentation is consistent with the observed variability in activation patching results, suggesting that circuit-level interpretations may differ notably even for similar inputs.
@@ 8 Conclusion
We present TinySQL, a dataset with increasingly complex SQL subsets and models of varying sizes. While not a replacement for real-world data, it supports controlled circuit discovery. Our analysis of EAP, SAE, and Logit Lens reveals both strengths and limitations, offering a testbed for advancing interpretability methods.
@@
While our approach provides meaningful insights into SQL generation mechanisms, several limitations remain. Interpretability results are sensitive to the choice of method and ablation strategy: zero-ablation offers robustness but inflates circuit size, while mean-ablation yields smaller circuits with unclear causal impact. Our EAP technique scales less effectively to larger models, and even the best-performing circuits recover only partial accuracy, suggesting that key mechanisms may remain undetected. We applied SAEs to CS4/5 to explore shared structure, but did not apply EAP due to its computational overhead and diminishing returns in deeper tasks, which require more targeted and scalable methods.
@@ C.1 Dataset Statistics for CSx_Syn
C.2 Dataset Generation Methodology: CSx, CSx_Syn
Step 2: Target Query Generation. Based on command set level, we generate SQL queries with increasing complexity. CS1 creates basic SELECT-FROM statements, selecting random schema columns. CS2 adds ORDER BY clauses (90% probability), randomly choosing ascending or descending order. CS3 introduces aggregate functions (COUNT, SUM, AVG, MIN, MAX), applied to numeric columns with an 85% probability, ensuring compatibility with data types (e.g., SUM and AVG for numeric fields). CS4 adds WHERE clauses (80% probability) with 1–3 type-consistent filters per query. CS5 introduces a second table with a JOIN clause on a shared field when schema compatibility allows.
C.3 Dataset Generation Methodology: CSx_Nat
We additionally explore a synthetic dataset generation method using Gretel AI’s Data Designer [^1] https://gretel.ai/navigator/data-designer. By defining two categorical seed columns (table_name and instruction_phrase) and several generation columns (e.g., column_names, selected_columns), we prompt Gretel’s language models to produce varied table names, column sets, and natural instructions. We then assemble them into structured SQL queries, applying post-processing rules for syntactic correctness. This approach further diversifies our text-to-SQL dataset by introducing additional schema and instruction variety.
@@ C.3 Dataset Generation Methodology: CSx_Nat
C.4 Dataset Examples
Table 6 shows representative examples from each category (CS1–CS5) in the TinySQL dataset. Each example includes a natural language question, a SQL statement and the corresponding SQL query.
Dataset Create Statement English Prompt SQL Statement
CS1 CREATE TABLE surveys ( file_type VARCHAR(100), expires_at DATETIME, comment TEXT, message TEXT, birthday DATE ) Get a readout of expires_at, birthday, comment, message and file_type from surveys SELECT expires_at, birthday, comment, message, file_type FROM surveys
CS1-SYN CREATE TABLE search_filters ( downloads BIGINT, passed BOOLEAN, website VARCHAR(255), discount DECIMAL(10,2), birthday DATE, is_active BOOLEAN, sequence SMALLINT, last_message TEXT, meta_keywords VARCHAR(500), response_id INTEGER, reference TEXT ) Can you get me recent message, date of birth, reply id, meta_keywords, url, downloads and referral from refinement options? SELECT last_message, birthday, response_id, meta_keywords, website, downloads, reference FROM search_filters
CS2 CREATE TABLE terms ( vote_id INTEGER, start_date DATE, mac_address VARCHAR(17), last_message TEXT, signature BLOB, middle_name VARCHAR(100), coordinates POINT, address VARCHAR(255), meta_description TEXT ) Can you get me coordinates, signature, address, start_date and mac_address from terms? top coordinates, best meta_description, with the highest start_date, starting with the highest last_message, starting with the highest mac_address SELECT coordinates, signature, address, start_date, mac_address FROM terms ORDER BY coordinates DESC, meta_description DESC, start_date DESC, last_message DESC, mac_address DESC
CS2-SYN CREATE TABLE configurations ( subject TEXT, image TEXT, response TEXT, major VARCHAR(100), company VARCHAR(255), session_id VARCHAR(100), city VARCHAR(100), min_score INTEGER, actual_time INTEGER, manufacturer VARCHAR(255) ) Looking in setup details, show me major, real duration, title, city, maker, session_id, photo, minimum points and organization worst company SELECT major, actual_time, subject, city, manufacturer, session_id, image, min_score, company FROM configurations ORDER BY company ASC
CS3 CREATE TABLE conversions ( total_price DECIMAL(10,2), approved_at TIMESTAMP, content LONGTEXT, estimated_time SMALLINT, permissions TEXT, first_name VARCHAR(100), salt VARCHAR(32), full_name VARCHAR(150) ) Read out total count total_price, how many full_name, times approved_at, tally salt, content, instances of permissions, first_name and how many estimated_time from conversions arranged by first_name, oldest permissions, structured by estimated_time SELECT COUNT(total_price) AS COUNT_total_price, COUNT(full_name) AS COUNT_full_name, COUNT(approved_at) AS COUNT_approved_at, COUNT(salt) AS COUNT_salt, content, COUNT(permissions) AS COUNT_permissions, first_name, COUNT(estimated_time) AS COUNT_estimated_time FROM conversions ORDER BY first_name ASC, permissions ASC, estimated_time ASC
CS3-SYN CREATE TABLE product_versions ( skills TEXT, location POINT ) Can you get me latest skills and location from releases? sorted alphabetically by location SELECT MAX(skills) AS MAX_skills, location FROM product_versions ORDER BY location ASC
CS4 CREATE TABLE data ( max_score INTEGER, gender CHAR(1), region GEOMETRY, location GEOMETRY, middle_name VARCHAR(100), salt VARCHAR(32), last_message TEXT ) Bring up middle_name, last_message, location, salt, region, max_score and gender from information where salt is containing ’%z%’ priority ordered by max_score, in reverse alphabetical order of location SELECT middle_name, last_message, location, salt, region, max_score, gender FROM data WHERE salt LIKE ’%z%’ ORDER BY max_score DESC, location DESC
CS5 CREATE TABLE workflow_templates ( like_id INTEGER, heading DECIMAL(5,2), is_verified BOOLEAN, share_id BIGINT, average_score FLOAT, rating DECIMAL(3,2), time TIME, certification VARCHAR(255), name VARCHAR(100), city VARCHAR(100) ) I need a list of verified status, like_id, time, average_score, designation, city, heading, rating, certification and share_id from procedure patterns join with pairs on like_id equals user_id ordered numerically by average_score, ordered by date of is_verified, alphabetically by certification, grouped by rating SELECT is_verified, like_id, time, average_score, name, city, heading, rating, certification, share_id FROM workflow_templates JOIN pairs ON workflow_templates.like_id = pairs.user_id ORDER BY average_score ASC, is_verified ASC, certification ASC, rating ASC
Table 6: Examples from TinySQL dataset across five compositional subsets (CS1–CS5) and CS1-SYN, CS2-SYN, CS3-SYN data.
@@ Appendix D Trained Models
Model Architectures.
The three models vary in size and architecture. BM1 (TinyStories-33M) has 2 transformer layers and 16 attention heads. BM2 (Qwen2.5-0.5B-Instruct) is composed of 24 layers with 14 attention heads. BM3 (Llama-3.2-1B-Instruct) consists of 16 layers and 32 attention heads. These configurations reflect each model’s respective scale and capacity.
@@
Appendix E Data set entry characteristics
For brevity, the main text of this paper uses short examples. The dataset entries are considerably longer. These are some characteristics of the generated datasets:
  • All CREATE TABLE clause contains (randomly) 2 to 12 columns.
  • All SELECT clause contains (randomly) 1 to the number of table columns.
  • In data set CS2, the ORDER BY clause is added and contains (randomly) 1 to the number of table columns. The ASC or DESC ordering is chosen randomly. An ORDER BY clause is added to 90% of entries.
  • In data set CS3, each SELECT column is randomly assigned a valid (depending on the column data type) aggregate from the list: SUM, AVG, MIN, MAX, COUNT and “”. That is, 2̃0% of columns have no aggregate.
  • In data set CS4, we extend CS3 by adding a type-aware WHERE clause to 80% of queries, selecting 1–3 distinct columns and type appropriate operators/values.
  • In data set CS5, we introduce a second table and a JOIN clause that matches columns with the same base type. If there are no columns in the second table that share a common type with the table already in the statement, we do not add a JOIN clause. This results in 73.0% of the train split for CS5 containing a JOIN.
Table 8 shows the maximum length of entries in each dataset.
Dataset English Prompt Create statement SQL statement
CS1 19 words, 169 chars 29 words, 299 chars 15 words, 159 chars
CS2 63 words, 476 chars 29 words, 308 chars 40 words, 325 chars
CS3 78 words, 579 chars 29 words, 296 chars 55 words, 564 chars
Table 8: Lengths of the longest entries in words and characters per dataset. For the synonym datasets, the English Prompts are 1̃0% longer.
Dataset Rarity Lexical Density Readab- ility
CS1_Synonyms 0.602 0.580 45.533
CS2_Synonyms 0.525 0.565 29.378
CS3_Synonyms 0.449 0.608 29.850
CS4_Synonyms 0.413 0.579 28.722
CS5_Synonyms 0.443 0.589 25.742
Table 9: Linguistic characteristics of CS*_Synonyms datasets.
Appendix F Activation Patching Results
Across prompts across all classes, we find that the activation patching results differ quite significantly from each other, even for prompts in the same class. We provide a sample of 4 examples in Figure 8, and we note that this variation persists throughout all examples.
@@ Appendix F Activation Patching Results
Appendix G EAP Circuits Identification
G.1 Edge Attribution Patching (EAP) Results
@@
G.1.1 Features used per dataset
@@ G.1.2 EAP results
G.2 Ablation Study Results
@@
G.3 Multi-Layer Perceptron (MLP) Ablations
@@
G.4 Full-Layer Ablations
@@ Appendix H SAE analysis
We show now the SAE features most involved in filling out the table name tokens, in particular for BM1_CS1_Syn, BM1_CS2_Syn and BM1_CS3_Syn. See Figures 9, 10 and 11 for the top activating SAE features by average magnitude.
@@
We can map these to the actual attention outputs, see for instance Figures 12, 12 and 12. We average over these to determine the influential attention heads.
@@
Figure 12: Attention outputs by head for selected SAE features across models and layers. Top row: BM1 with two different contexts; Bottom: BM3 highlighting deeper-layer behavior.
Appendix I SAE Circuits Results
We present additional results from our Synthetic Activation Experiments (SAEs) highlighting the identified circuit components and their recovered accuracy after zero-ablation across different model and dataset configurations (Table 13).
Model Dataset Heads and MLPs Recovered Accuracy
Layer 0 Layer 1
BM1_CS3_Syn CS2 13, 1, 8, 15, 10, 12, 14, 3, 7 81.41%
All MLP neurons All MLP neurons
BM1_CS4_Syn CS4 8, 1, 10, 13, 3, 6, 12, 9, 2, 15 82.19%
750 MLP neurons
BM1_CS5_Syn CS5 8, 10, 1, 13, 3, 12, 6, 9, 2, 7, 4 70.00%
All MLP neurions
Table 13: Additional SAE-based circuit results for BM1-CS3→CS2, CS4, and CS5 configurations. Recovered accuracy reflects performance after zero-ablation of all model components not selected by high-AUC SAE features.
Table.14 details the layer-wise circuits identified by Synthetic Activation Experiments (SAEs) for the BM2_CS1_Syn model, demonstrating more granular selection of individual attention heads compared to the broader patterns found by EAP.
Model Layer Selected Heads Recovered Accuracy
BM2_CS1_Syn L0 12, 7, 6, 0, 10, 2, 1, 9, 4
L1 All Heads
L2 0, 11, 4, 5, 9, 8, 13, 1, 2
L3 0, 11, 10, 12, 3, 1, 9, 4
L4 All Heads
L5 All Heads
L6 5, 0, 2, 13, 6, 8, 4, 12, 7, 9
L7 9, 1, 6, 0, 2, 4, 11, 8, 3
L8 All Heads
L9 0, 3, 8, 4, 13, 9, 7, 10, 1
L10 All Heads
L11 All Heads 88.1%
L12 0, 8, 7, 11, 3, 1, 5, 13
L13 0, 3, 11, 7, 1, 2, 4, 10, 9
L14 All Heads
L15 9, 7, 10, 1, 4, 3, 8, 13, 5, 0
L16 All Heads
L17 13, 5, 4, 10, 7, 0, 8, 11
L18 3, 0, 5, 13, 11, 8, 4, 9, 1, 6
L19 3, 0, 7, 4, 11, 10, 13, 1, 5, 12
L20 0, 3, 7, 2, 1, 13, 10
L21 0, 3, 7, 1, 10, 6, 11, 9
L22 0, 3, 7, 1, 2, 11, 10, 6
L23 12, 1, 9, 13, 5, 2, 11
Table 14: Layer-wise SAE-selected circuits for BM2_CS1_Syn. Compared to EAP, the circuits identified by SAEs exhibit finer granularity, isolating individual attention heads rather than entire layers.
Appendix J Mean-Ablation Circuits
Circuits in Small 2-Layer Models.
We begin by analyzing models with only two transformer layers (e.g., BM1 and BM3 variants). The mean ablation results for these models reveal small, interpretable sets of attention heads responsible for accurate task performance. Table 15 summarizes the critical heads involved in each model-task pair and their corresponding recovered accuracy after ablation.
Model Dataset Important Heads (Layer1) Recovered Accuracy Component %
BM1-CS1 CS1 11, 3, 1, 8 93.68% 12.5%
BM1-CS3 CS1 11, 14, 3, 8, 2 93.8% 15.62%
BM1-CS3 CS2 11 99.0% 3.12%
BM1-CS3 CS3 11, 14, 3, 0, 13 73.43% 15.62%
BM3-CS1 CS1 0, 5 99.2% 8.3%
Table 15: Mean Ablation Results for Attention Circuits in BM1 and BM3 Models
Circuits in a Deeper Model: BM2-CS1.
To investigate whether similar sparse circuits exist in deeper models, we perform a layer-wise mean ablation analysis on BM2-CS1, which has 24 layers. For each layer, we identify key heads contributing to recovery and report both the area under the curve (AUC) and the recovered accuracy. Table 16 presents these results, showing how different layers contribute to robust model behavior via distributed yet sparse mechanisms.
Layer Key Heads Recovered Accuracy
0 12 98.6%
1 7 98.4%
2 0 98.2%
3 0, 11, 10, 12, 3, 1, 9, 4, 13 98.3%
4 2 98.3%
5 0 98.4%
6 5 97.9%
7 9 98.2%
8 11 98.2%
9 3 98.39%
10 0 98.2%
11 3 97.8%
12 7 98.2%
13 0 98.2%
14 0 97.9%
15 10, 7, 4, 13, 1, 9, 3, 8
16 0, 7, 13, 10, 3, 11, 2, 12, 1 97.9%
17 13, 0, 5, 4, 1, 3, 10 98.2%
18 3, 4, 5, 11, 0, 13, 12, 9, 6, 8 98.2%
19 3, 0, 13, 10, 7, 12, 5, 4, 1 97.9%
20 0, 3, 7, 2, 6, 4, 10, 1, 8 97.9%
21 7, 13, 1, 6, 3, 8, 10, 9, 0 98.2%
22 0, 3, 1, 6, 7, 8, 9, 12, 10 97.9%
23 12, 13, 1, 6, 9, 11, 5, 0, 2, 7 98.8%
Table 16: Mean Ablation Results Across Layers for BM2-CS1
Efficiency of Model Component Usage.
Beyond identifying which layers and heads are important, we quantify how much of the model is actually needed for successful task completion. Table 17 reports the percentage of total components (e.g., attention heads or MLP neurons) that were retained in each ablation configuration while still achieving high performance. These results indicate that high accuracy can often be recovered using only a small fraction of the model’s total capacity.
Model MI Method Component Type Percentage of Model Components
BM1-CS1-SYN CS1 EAP 12.5%
BM1-CS3-SYN CS1 EAP 15.62%
BM1-CS3-SYN CS2 EAP 3.12%
BM1-CS3-SYN CS3 EAP 15.62%
BM3-CS1-SYN CS1 EAP 8.3%
BM2-CS1-SYN CS1 EAP 12.5%
BM1-CS1-SYN CS1 SAE 17.26%
BM1-CS3-SYN CS1 SAE 22.5%
BM1-CS3-SYN CS2 SAE 18.75%
BM1-CS3-SYN CS3 SAE 28.97%
BM2-CS1-SYN CS1 SAE 60.71%
Table 17: Percentage of Model Components Used in Successful Recovery. In EAP, we only consider the percentage of Attention Heads as we only had access to those.
Appendix K Comparison of EAP-Selected Circuits and Random Circuits
Table 18 presents a comparison between the recovered accuracies of circuits selected by Edge Attribution Patching (EAP) and the average accuracies of randomly selected circuits of comparable size across different model and dataset configurations.
Model EAP Circuit Accuracy Average Random Circuit Accuracy
BM1-CS1-SYN $\to$ CS1 85.4% 54.55%
BM1-CS3-SYN $\to$ CS1 97.6% 82.76%
BM1-CS3-SYN $\to$ CS2 74.6% 57.83%
BM1-CS3-SYN $\to$ CS3 78.3% 66.66%
BM2-CS1-SYN $\to$ CS1 71.6% 30.0%
Table 18: Recovered accuracy comparison between EAP-selected circuits and randomly selected circuits of similar size.
Appendix L Additional LogitLens Visualizations
To further support our interpretation of two-phase SQL query generation, we include additional LogitLens visualizations across other benchmarks and model configurations.
In each case, we consistently observe the same dynamics: an early spike in the logit probabilities of SQL-specific keywords, followed by a later emergence of table name tokens, indicating a transition from syntactic intent planning to context-aware content grounding.
These findings generalize across schema types and query contexts, reinforcing the idea that language models internally decompose structured generation tasks into distinct stages aligned with human-interpretable reasoning processes.
(a) BM3-CS3
(b) BM3-CS1
Figure 13: Additional LogitLens visualizations across six benchmarks, illustrating consistent two-phase behavior: early SQL intent formation followed by table resolution.
@@ Appendix I SAE Circuits Results
Figure 17: Ablation results showing accuracy trends as more attention heads remain unablated. (Top-left) BM1_CS1_Syn with CS1 data, (Top-right) BM1_CS3_Syn with CS1 data, (Middle-left) BM1_CS3_Syn with CS2 data, (Middle-right) BM1_CS3_Syn with CS3 data, (Bottom) BM2_CS1_Syn with CS1 data.